40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 10 01:10:56 2018, gautam, Update, LSC, Some vertex locking restored PRMIcarrier.png
    Reply  Tue Aug 14 00:27:55 2018, gautam, Update, LSC, Locking prep 
       Reply  Tue Aug 14 02:01:12 2018, gautam, Update, LSC, DRMI locking - partial success DRMI.png
Message ID: 14160     Entry time: Tue Aug 14 00:27:55 2018     In reply to: 14152     Reply to this: 14162
Author: gautam 
Type: Update 
Category: LSC 
Subject: Locking prep 

In preparation for attempting some DRMI locking, I did the following:

  • Slow machine reboots for unresponsive c1psl, c1susaux and c1iscaux. The latter requried a manual burtrestore to recover the usual LSC PD whitening settings.
  • Shuttered AUX laser (which was on Standby anyways) - we should really install a remotely controllable shutter for this on the AS table.
  • Re-aligned PMC (half turn of knob in yaw, full turn in pitch) - IMC transmission 15,000cts ---> 15,600cts.
  • Squished sat. box cables at ITMX and ETMX.

Not related to this work, but I turned the Agilent NA off since we aren't using it immediately.

ELOG V3.1.3-